Simplify 10^(1/3)n ×15^(1/2)n ×6^(1/6)n ÷45^(1/3)n
1
Expert's answer
2020-07-07T18:45:31-0400

10(1/3)n × 15(1/2)n × 6(1/6)n ÷ 45(1/3)n

Get the factors of:

10=2×5

15=3×5

6=2×3

45=32×5

{2(1/3)n × 5(1/3)n × 3(1/2)n × 5(1/2)n × 2(1/6)n × 3(1/6)n} ÷{3(2/3)n × 5(1/3)n}

Group common bases

{2(1/3)n × 2(1/6)n} × {3(1/2)n × 3(1/6)n ÷ 3(2/3)n} × {5(1/2)n × 5(1/3)n ÷ 5(1/3)n}

2(1/2)n × 30 × 5(1/2)n

10(1/2)n